# Evaluating Memory Capability in Continuous Lifelog Scenario

### 13 Apr 2026 v1

@@ Abstract
Nowadays, wearable devices can continuously lifelog ambient conversations, creating substantial opportunities for memory systems. However, existing benchmarks primarily focus on online one-on-one chatting or human-AI interactions, thus neglecting the unique demands of real-world scenarios. Given the scarcity of public lifelogging audio datasets, we propose a hierarchical synthesis framework to curate LifeDialBench, a novel benchmark comprising two complementary subsets: EgoMem, built on real-world egocentric videos, and LifeMem, constructed using simulated virtual community. Crucially, to address the issue of temporal leakage in traditional offline settings, we propose an Online Evaluation protocol that strictly adheres to temporal causality, ensuring systems are evaluated in a realistic streaming fashion. Our experimental results reveal a counterintuitive finding: current sophisticated memory systems fail to outperform a simple RAG-based baseline. This highlights the detrimental impact of over-designed structures and lossy compression in current approaches, emphasizing the necessity of high-fidelity context preservation for lifelog scenarios. We release our code and data at https://github.com/qys77714/LifeDialBench.
@@ 1 Introduction
We evaluate four representative memory systems on LifeDialBench. Unexpectedly, a simple RAG baseline consistently outperforms specialized methods, highlighting the advantage of raw text preservation over the lossy compression inherent in current designs. Furthermore, we identify temporal retrieval as a universal bottleneck across all methods. Crucially, our analysis validates the necessity of online evaluation, demonstrating that traditional offline settings distort assessment by permitting temporal leakage and misjudging dynamic memory updates. These findings collectively underscore the dual importance of context fidelity and strict temporal linearity in next-generation lifelog systems.
@@ 3 Benchmarking Lifelog Memory
To explore the long-term memory capacity of agents in continuous dialogue lifelogs, we specifically construct a benchmark named LifeDialBench with two complementary subsets for egocentric memory(Cheng et al., 2024), as illustrated in Figure˜1. The first subset, named EgoMem, is constructed based on the existing EgoLife dataset (Yang et al., 2025b) which contains daily video recording across seven days. Moreover, to mimic the continuous dialogue lifelogs in real-life with more time span and scene diversity, we further adopt data synthesis to construct a year-long subset, named LifeMem. Both subsets are constructed with a hierarchical life simulation framework, where we use bottom-up and top-down manners due to different data sources. Specifically, all dialogue and summary content are generated using Qwen3-235B-Instruct111corresponds to Qwen3-235B-A22B-Instruct-2507.
@@ 3.2 EgoMem
To ensure our synthesized data reflects the natural dynamics of daily communication rather than overly formalized written prose, we implement a human-LLM collaborative review pipeline (Yuksekgonul et al., 2024). Beyond standard Factual Consistency checks against the source summaries, we specifically optimize for Conversational Naturalness. Human annotators utilize LLM-flagged feedback to iteratively revise the text, ensuring the dialogues capture casual interactions, appropriate tone, and implicit contexts typical of real-world continuous recordings. Comprehensive details regarding ASR limitations, granularity selection, and the full review workflow are provided in Appendix C.
@@ 3.4 Question-Answering Pairs Curation
We prompt Qwen3-235B-Thinking 222corresponds to Qwen3-235B-A22B-Thinking-2507  (Yang et al., 2025a) to synthesize QA pairs spanning multiple temporal granularities. To ensure rigorous evaluation, we implement a strict quality assurance pipeline.
@@
3) Answerability Check: We prompt qwen3-max 333corresponds to qwen3-max-2025-09-23 to answer the generated questions. To mitigate random guessing (a 25% chance in MCQ), the evaluator must output explicit evidence rationales before selecting an option. Unanswerable questions are discarded, and correct ones undergo human random sampling to mitigate LLM self-preference bias.
Section˜3.4 details the filtering statistics and near-perfect evaluator accuracy for LifeMem, confirming the genuine answerability of the curated queries. Ultimately, the benchmark retains 1,774 questions for EgoMem and 1,717 for LifeMem. Detailed prompts are in Appendix F.3.
@@ 3.5 Dataset Statistics
For QA pairs, Figure˜6 in section˜G.2 shows the proportion of each question type after filtering. The four types—event content recall, event detail retrieval, multi-hop event reasoning, and temporal information QA—are distributed as 25.3/25.0/25.0/24.6 in LifeMem and 25.1/25.1/24.9/24.9 in EgoMem, ensuring a balanced evaluation of memory systems across multifaceted retrieval and reasoning capabilities.
@@ 4.1 Experimental Setup
To assess the performance of current mainstream memory systems in the context of continuous dialogue lifelogs and to derive insights, we select four representative memory systems for evaluation on LifeDialBench: (1) RAG (Lewis et al., 2021): A straightforward retrieval-augmented generation (RAG) baseline that directly stores and retrieves text chunks; (2) A-Mem (Xu et al., 2025): An enhanced variant of RAG that augments the retrieval process with additional semantic signals — such as context, tags, keywords, and links — to improve the representation and utilization of stored text; (3) MemOS (Li et al., 2025): A memory system that does not retain raw input, but instead abstracts each input segment into a single structured representation—including summaries, titles, and semantic tags; and (4) Mem0 (Chhikara et al., 2025): A memory paradigm that similarly discards raw input, but extracts and stores multiple concise factual statements per input segment. More information about the implementation details of these methods can be found in Appendix˜H.
Experiments on EgoMem and LifeMem are conducted in an online setting using GPT-4o-mini 444corresponds to GPT-4o-mini and qwen-plus 555corresponds to qwen-plus-2025-12-01
@@ 6 Conclusion
Our extensive experiments yield a counter-intuitive yet critical finding: despite the growing complexity of agentic memory architectures (e.g., graph-based or summary-based systems), simple raw-text retrieval (RAG) consistently outperforms structured memory paradigms. Our analysis reveals that current abstraction mechanisms introduce "lossy compression," stripping away vital contextual details required for precise grounding in continuous streams. Furthermore, we demonstrate that offline metrics suffer from severe future context contamination, rendering them unreliable. By establishing these insights, LifeDialBench serves not only as a testbed but as a call to action for the community to rethink memory design—shifting focus from aggressive abstraction back to high-fidelity context preservation and efficient temporal retrieval.
7 Ethics Statement
@@ Appendix B Details of human-in-the-loop review
Appendix C Details of EgoMem Construction
@@
C.1 Limitations of Direct ASR Transcription
@@
C.2 Justification for Granularity and Context Window
@@
C.3 Iterative Review and Conversational Grounding
@@
Appendix D Sensitivity Analysis
D.1 Impact of Backbone Capability
@@
Appendix E Benchmark Data Samples
E.1 Examples of dataset
@@
E.2 Question Types and Examples
@@
Appendix F Prompts
F.1 Judge Prompt used for Open-Ended Format
@@
F.2 Lifelog Generation Prompt
@@
F.3 Question Generation Prompt
@@
Appendix G Additional Results
G.1 Offline Evaluation Results
@@
G.2 Question Types Distribution
@@
Appendix H Memory Systems
H.1 Description of Memory Systems

### 17 Apr 2026 v2

@@ Abstract
Nowadays, wearable devices can continuously lifelog ambient conversations, creating substantial opportunities for memory systems. However, existing benchmarks primarily focus on online one-on-one chatting or human-AI interactions, thus neglecting the unique demands of real-world scenarios. Given the scarcity of public lifelogging audio datasets, we propose a hierarchical synthesis framework to curate LifeDialBench, a novel benchmark comprising two complementary subsets: EgoMem, built on real-world egocentric videos, and LifeMem, constructed using simulated virtual community. Crucially, to address the issue of temporal leakage in traditional offline settings, we propose an Online Evaluation protocol that strictly adheres to temporal causality, ensuring systems are evaluated in a realistic streaming fashion. Our experimental results reveal a counterintuitive finding: current sophisticated memory systems fail to outperform a simple RAG-based baseline. This highlights the detrimental impact of over-designed structures and lossy compression in current approaches, emphasizing the necessity of high-fidelity context preservation for lifelog scenarios.
@@ 1 Introduction
We evaluate four representative memory systems on LifeDialBench. Unexpectedly, a simple RAG baseline consistently outperforms specialized methods, highlighting the advantage of raw text preservation over the lossy compression inherent in current designs. Furthermore, we identify temporal retrieval as a universal bottleneck across all methods. Crucially, our analysis validates the necessity of online evaluation, demonstrating that traditional offline settings distort assessment by permitting temporal leakage and misjudging dynamic memory updates. These findings collectively underscore the dual importance of context fidelity and strict temporal linearity in next-generation lifelog systems.  [^1] We release our code and data at https://github.com/RayNeo-AI-2025/LifeDialBench.
@@ 3 Benchmarking Lifelog Memory
To explore the long-term memory capacity of agents in continuous dialogue lifelogs, we specifically construct a benchmark named LifeDialBench with two complementary subsets for egocentric memory(Cheng et al., 2024), as illustrated in Figure˜1. The first subset, named EgoMem, is constructed based on the existing EgoLife dataset (Yang et al., 2025b) which contains daily video recording across seven days. Moreover, to mimic the continuous dialogue lifelogs in real-life with more time span and scene diversity, we further adopt data synthesis to construct a year-long subset, named LifeMem. Both subsets are constructed with a hierarchical life simulation framework, where we use bottom-up and top-down manners due to different data sources. Specifically, all dialogue and summary content are generated using Qwen3-235B-Instruct [^2] Corresponds to Qwen3-235B-A22B-Instruct-2507.
@@ 3.2 EgoMem
To ensure our synthesized data reflects the natural dynamics of daily communication rather than overly formalized written prose, we implement a human-LLM collaborative review pipeline (Yuksekgonul et al., 2024). Beyond standard Factual Consistency checks against the source summaries, we specifically optimize for Conversational Naturalness. Human annotators utilize LLM-flagged feedback to iteratively revise the text, ensuring the dialogues capture casual interactions, appropriate tone, and implicit contexts typical of real-world continuous recordings. Comprehensive details regarding ASR limitations, granularity selection, and the full review workflow are provided in Appendix D.
@@ 3.4 Question-Answering Pairs Curation
We prompt Qwen3-235B-Thinking  [^3] Corresponds to Qwen3-235B-A22B-Thinking-2507  (Yang et al., 2025a) to synthesize QA pairs spanning multiple temporal granularities. To ensure rigorous evaluation, we implement a strict quality assurance pipeline.
@@
3) Answerability Check: We prompt qwen3-max  [^4] Corresponds to qwen3-max-2025-09-23 to answer the generated questions. To mitigate random guessing (a 25% chance in MCQ), the evaluator must output explicit evidence rationales before selecting an option. Unanswerable questions are discarded, and correct ones undergo human random sampling to mitigate LLM self-preference bias.
Section˜3.4 details the filtering statistics and near-perfect evaluator accuracy for LifeMem, confirming the genuine answerability of the curated queries. Ultimately, the benchmark retains 1,774 questions for EgoMem and 1,717 for LifeMem. Detailed prompts are in Appendix G.3.
@@ 3.5 Dataset Statistics
For QA pairs, Figure˜6 in section˜H.2 shows the proportion of each question type after filtering. The four types—event content recall, event detail retrieval, multi-hop event reasoning, and temporal information QA—are distributed as 25.3/25.0/25.0/24.6 in LifeMem and 25.1/25.1/24.9/24.9 in EgoMem, ensuring a balanced evaluation of memory systems across multifaceted retrieval and reasoning capabilities.
@@ 4.1 Experimental Setup
To assess the performance of current mainstream memory systems in the context of continuous dialogue lifelogs and to derive insights, we select four representative memory systems for evaluation on LifeDialBench: (1) RAG (Lewis et al., 2021): A straightforward retrieval-augmented generation (RAG) baseline that directly stores and retrieves text chunks; (2) A-Mem (Xu et al., 2025): An enhanced variant of RAG that augments the retrieval process with additional semantic signals — such as context, tags, keywords, and links — to improve the representation and utilization of stored text; (3) MemOS (Li et al., 2025): A memory system that does not retain raw input, but instead abstracts each input segment into a single structured representation—including summaries, titles, and semantic tags; and (4) Mem0 (Chhikara et al., 2025): A memory paradigm that similarly discards raw input, but extracts and stores multiple concise factual statements per input segment. More information about the implementation details of these methods can be found in Appendix˜I.
Experiments on EgoMem and LifeMem are conducted in an online setting using GPT-4o-mini  [^5] Corresponds to GPT-4o-mini and qwen-plus  [^6] Corresponds to qwen-plus-2025-12-01
@@ 6 Conclusion
Our extensive experiments yield a counter-intuitive yet critical finding: despite the growing complexity of agentic memory architectures (e.g., graph-based or summary-based systems), simple raw-text retrieval (RAG) consistently outperforms structured memory paradigms. Our analysis reveals that current abstraction mechanisms introduce "lossy compression," stripping away vital contextual details required for precise grounding in continuous streams. Furthermore, we demonstrate that offline metrics suffer from severe future context contamination, rendering them unreliable. By establishing these insights, LifeDialBench serves not only as a testbed but as a call to action for the community to rethink memory design—shifting focus from aggressive abstraction back to high-fidelity context preservation.
Ethics Statement
@@ Appendix B Details of human-in-the-loop review
Appendix C Annotator Details
All human annotators are professional data annotation employees of RayNeo AI, recruited internally from the company’s dedicated annotation team. Compensation was provided in accordance with the company’s standard salary structure, commensurate with their professional roles.
Appendix D Details of EgoMem Construction
@@
D.1 Limitations of Direct ASR Transcription
@@
D.2 Justification for Granularity and Context Window
@@
D.3 Iterative Review and Conversational Grounding
@@
Appendix E Sensitivity Analysis
E.1 Impact of Backbone Capability
@@
Appendix F Benchmark Data Samples
F.1 Examples of dataset
@@
F.2 Question Types and Examples
@@
Appendix G Prompts
G.1 Judge Prompt used for Open-Ended Format
@@
G.2 Lifelog Generation Prompt
@@
G.3 Question Generation Prompt
@@
Appendix H Additional Results
H.1 Offline Evaluation Results
@@
H.2 Question Types Distribution
@@
Appendix I Memory Systems
I.1 Description of Memory Systems